# Evaluating small vision-language models as AI assistants for radio astronomical source analysis tasks

### v2

@@ A.2 Fine-grained radio datasets
The dataset currently contains 12,774 annotated radio images taken from different surveys, such as the VLA FIRST (Becker et al., 1995), ATCA Scorpio (Umana et al., 2015), and ASKAP-EMU Scorpio (Umana et al., 2021) surveys. The annotation data include bounding boxes, segmentation masks and classification labels for all radio object identified in the images (38,342 objects, including both real and spurious sources). Objects are classified into five possible classes: SPURIOUS, COMPACT, EXTENDED, EXTENDED-MULTISLAND, FLAGGED. A detailed explanation of the labelling scheme is provided in the reference publication (Riggi et al., 2023). The entire dataset was used to produce the Q&A training dataset.
@@
The dataset currently contains 10,414 annotated radio images taken from the ASKAP EMU pilot survey (Norris et al., 2021), each containing at least one extended radio source. Annotation data have been extracted from EMU pilot RG-CAT catalogue (Gupta et al, 2024), including bounding boxes and classification labels for radio objects present in the images. Objects in the original catalogue are classified into six possible radio galaxy morphology classes:
@@
From a visual inspection of the data, we note that various objects classified as compact (C) should be rather considered as belonging to the EXTENDED class in the classification scheme adopted in the caesar-mrcnn dataset (see previous paragraph). To make the two fine-grained datasets more comparable, we applied the following processing steps. We first extracted objects from EMU pilot images using caesar-mrcnn trained model (Riggi et al., 2023). As a result, we obtained a list of detected objects classified with the caesar-mrcnn classification scheme, that was cross-matched with the original RG-CAT object collection. This was extended and complemented according to the match results. Objects with a match ($\sim$78%) were also given a caesar-mrcnn label. Objects detected by the caesar-mrcnn model but missed in the RG-CAT were added to the final collection, including a total of 231,439 objects. The obtained source annotations were taken into account to generate the conversational dataset.
@@ Appendix B Image multi-modal benchmarks
This benchmark [^13] https://huggingface.co/datasets/lmms-lab/ai2d consists of 3,088 image-based Q&A pairs on annotated grade school science diagrams from the AI2 Diagrams (AI2D) dataset (Kembhavi et al., 2016).
@@
This benchmark [^14] https://huggingface.co/datasets/lmms-lab/ChartQA contains 2,500 image-based Q&A pairs on real-world charts in various formats (pie, bar) from the ChartQA dataset (Masry et al., 2022).
@@
This benchmark [^15] https://huggingface.co/datasets/lmms-lab/DocVQA contains 16,626 image-based Q&A pairs on document of various types and content, sourced from the DocVQA dataset (Mathew et al., 2021).
@@
This benchmark [^16] https://huggingface.co/datasets/lmms-lab/DocVQA, see InfographicVQA validation data split contains 2,801 image-based Q&A pairs on document infographics of various types and content, sourced from the InfographicVQA dataset (Mathew et al., 2022).
@@
This benchmark [^17] https://huggingface.co/datasets/lmms-lab/MME consists of 2,374 image-based Q&A pairs from the MME dataset (Fu et al., 2024), designed to evaluate multimodal models’ perception and cognition abilities. Perception tasks include OCR, recognition of coarse-grained objects (e.g., object presence, count, position, and color) and fine-grained objects (e.g., identification of movie posters, celebrities, scenes, landmarks and artworks). Cognition tasks cover common sense reasoning, numerical calculation, text translation, and code reasoning.
@@
This benchmark [^18] https://huggingface.co/datasets/lmms-lab/MMMU, see validation data split consists of 900 image-based Q&A pairs from the MMMU dataset (Yue et al., 2024), designed to assess multimodal perception and reasoning abilities across various image formats, incuding charts, diagrams, maps, tables, music sheets, and chemical structures. The images are sourced from college exams, quizzes, and textbooks spanning six disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering.
@@
This benchmark [^19] https://huggingface.co/datasets/Lin-Chen/MMStar contains 1,500 image-based Q&A pairs from the MMStar dataset (Chen et al., 2024), designed to evaluate multimodal models across six core capabilities: Coarse Perception, Fine-grained Perception, Instance Reasoning, Logical Reasoning, Science & Technology, Mathematics.
@@
This benchmark [^20] https://huggingface.co/datasets/echo840/OCRBench consists of 1,000 image-based Q&A pairs from the OCRBench dataset (Liu et al., 2024c), designed to assess Optical Character Recognition (OCR) capabilities across various domains, including multilingual text, handwritten text, non-semantic text, and mathematical expression recognition.
@@
This benchmark [^21] https://huggingface.co/datasets/lmms-lab/SEED-Bench contains 17,990 image-based Q&A pairs from the SEED-Bench dataset (Li et al., 2023c), designed to evaluate multimodal model capabilities across various domains, including scene understanding, instance identity/attribute/location/counting, spatial relations, instance interaction, visual reasoning, text recognition, action recognition/prediction, procedure understanding.
@@
This benchmark [^22] https://huggingface.co/datasets/lmms-lab/ScienceQA, see ScienceQA-IMG test data split consists of 2,017 image-based Q&A pairs from the ScienceQA dataset (Lu et al., 2022), featuring multiple-choice questions spanning diverse topics across three subjects: natural science, social science, and language science.
@@ Appendix C Supplementary plots/tables
radio-llava (Q&A + capt.) The image shows a radio astronomical cutout of an extended radio galaxy candidate, characterized by its elongated structure and multiple bright spots. The image shows a radio astronomical cutout with two prominent sources. One source appears as a bright, compact object near the center of the image, while the other is a more extended source located towards the right side. The background exhibits a complex pattern of concentric rings, likely resulting from the imaging process.
@@
Figure C.2: Class-averaged classification F1-scores obtained with the radio-llava model on B1–B6 radio benchmarks, comparing fine-tuning on the Q&A training dataset (black solid histograms, labelled as "default") with model variants (v1-v6, colored histograms), fine-tuned on the same dataset using alternative parameters (see text). The dashed black histogram represents the standard model evaluated on B1-B6 benchmarks using an alternative prompt.

### v3

@@ A.2 Fine-grained radio datasets
The dataset currently contains 12,774 annotated radio images taken from different surveys, such as the VLA FIRST Becker et al. (1995), ATCA Scorpio Umana et al. (2015), and ASKAP-EMU Scorpio Umana et al. (2021) surveys. The annotation data include bounding boxes, segmentation masks and classification labels for all radio object identified in the images (38,342 objects, including both real and spurious sources). Objects are classified into five possible classes: SPURIOUS, COMPACT, EXTENDED, EXTENDED-MULTISLAND, FLAGGED. A detailed explanation of the labelling scheme is provided in the reference publication Riggi et al. (2023). The entire dataset was used to produce the Q&A training dataset.
@@
The dataset currently contains 10,414 annotated radio images taken from the ASKAP EMU pilot survey Norris et al. (2021), each containing at least one extended radio source. Annotation data have been extracted from EMU pilot RG-CAT catalogue Gupta et al (2024), including bounding boxes and classification labels for radio objects present in the images. Objects in the original catalogue are classified into six possible radio galaxy morphology classes:
@@
From a visual inspection of the data, we note that various objects classified as compact (C) should be rather considered as belonging to the EXTENDED class in the classification scheme adopted in the caesar-mrcnn dataset (see previous paragraph). To make the two fine-grained datasets more comparable, we applied the following processing steps. We first extracted objects from EMU pilot images using caesar-mrcnn trained model Riggi et al. (2023). As a result, we obtained a list of detected objects classified with the caesar-mrcnn classification scheme, that was cross-matched with the original RG-CAT object collection. This was extended and complemented according to the match results. Objects with a match ($\sim$78%) were also given a caesar-mrcnn label. Objects detected by the caesar-mrcnn model but missed in the RG-CAT were added to the final collection, including a total of 231,439 objects. The obtained source annotations were taken into account to generate the conversational dataset.
@@ Appendix B Image multi-modal benchmarks
This benchmark [^14] https://huggingface.co/datasets/lmms-lab/ai2d consists of 3,088 image-based Q&A pairs on annotated grade school science diagrams from the AI2 Diagrams (AI2D) dataset Kembhavi et al. (2016).
@@
This benchmark [^15] https://huggingface.co/datasets/lmms-lab/ChartQA contains 2,500 image-based Q&A pairs on real-world charts in various formats (pie, bar) from the ChartQA dataset Masry et al. (2022).
@@
This benchmark [^16] https://huggingface.co/datasets/lmms-lab/DocVQA contains 16,626 image-based Q&A pairs on document of various types and content, sourced from the DocVQA dataset Mathew et al. (2021).
@@
This benchmark [^17] https://huggingface.co/datasets/lmms-lab/DocVQA, see InfographicVQA validation data split contains 2,801 image-based Q&A pairs on document infographics of various types and content, sourced from the InfographicVQA dataset Mathew et al. (2022).
@@
This benchmark [^18] https://huggingface.co/datasets/lmms-lab/MME consists of 2,374 image-based Q&A pairs from the MME dataset Fu et al. (2024), designed to evaluate multimodal models’ perception and cognition abilities. Perception tasks include OCR, recognition of coarse-grained objects (e.g., object presence, count, position, and colour) and fine-grained objects (e.g., identification of movie posters, celebrities, scenes, landmarks and artworks). Cognition tasks cover common sense reasoning, numerical calculation, text translation, and code reasoning.
@@
This benchmark [^19] https://huggingface.co/datasets/lmms-lab/MMMU, see validation data split consists of 900 image-based Q&A pairs from the MMMU dataset Yue et al. (2024), designed to assess multimodal perception and reasoning abilities across various image formats, including charts, diagrams, maps, tables, music sheets, and chemical structures. The images are sourced from college exams, quizzes, and textbooks spanning six disciplines: Art & Design, Business, Science, Health & Medicine, Humanities & Social Science, and Tech & Engineering.
@@
This benchmark [^20] https://huggingface.co/datasets/Lin-Chen/MMStar contains 1,500 image-based Q&A pairs from the MMStar dataset Chen et al. (2024), designed to evaluate multimodal models across six core capabilities: Coarse Perception, Fine-grained Perception, Instance Reasoning, Logical Reasoning, Science & Technology, Mathematics.
@@
This benchmark [^21] https://huggingface.co/datasets/echo840/OCRBench consists of 1,000 image-based Q&A pairs from the OCRBench dataset Liu et al. (2024c), designed to assess Optical Character Recognition (OCR) capabilities across various domains, including multilingual text, handwritten text, non-semantic text, and mathematical expression recognition.
@@
This benchmark [^22] https://huggingface.co/datasets/lmms-lab/SEED-Bench contains 17,990 image-based Q&A pairs from the SEED-Bench dataset Li et al. (2023c), designed to evaluate multimodal model capabilities across various domains, including scene understanding, instance identity/attribute/location/counting, spatial relations, instance interaction, visual reasoning, text recognition, action recognition/prediction, procedure understanding.
@@
This benchmark [^23] https://huggingface.co/datasets/lmms-lab/ScienceQA, see ScienceQA-IMG test data split consists of 2,017 image-based Q&A pairs from the ScienceQA dataset Lu et al. (2022), featuring multiple-choice questions spanning diverse topics across three subjects: natural science, social science, and language science.
@@ Appendix C Supplementary plots/tables
radio-llava (Q&A + capt.) The image shows a radio astronomical cutout of an extended radio galaxy candidate, characterised by its elongated structure and multiple bright spots. The image shows a radio astronomical cutout with two prominent sources. One source appears as a bright, compact object near the center of the image, while the other is a more extended source located towards the right side. The background exhibits a complex pattern of concentric rings, likely resulting from the imaging process.
@@
Figure C.2: Class-averaged classification F1-scores obtained with the radio-llava model on B1–B6 radio benchmarks, comparing fine-tuning on the Q&A training dataset (black solid histograms, labelled as "default") with model variants (v1-v6, coloured histograms), fine-tuned on the same dataset using alternative parameters (see text). The dashed black histogram represents the standard model evaluated on B1-B6 benchmarks using an alternative prompt.